# Cubic-based Prediction Approach for Large Volatility Matrix using High-Frequency Financial Data

### v1

@@ Abstract
In this paper, we develop a novel method for predicting future large volatility matrices based on high-dimensional factor-based Itô processes. Several studies have proposed volatility matrix prediction methods using parametric models to account for volatility dynamics. However, these methods often impose restrictions, such as constant eigenvectors over time. To generalize the factor structure, we construct a cubic (order-3 tensor) form of an integrated volatility matrix process, which can be decomposed into low-rank tensor and idiosyncratic tensor components. To predict conditional expected large volatility matrices, we introduce the Projected Tensor Principal Orthogonal componEnt Thresholding (PT-POET) procedure and establish its asymptotic properties. Finally, the advantages of PT-POET are also verified by a simulation study and illustrated by applying minimum variance portfolio allocation using high-frequency trading data.
@@ 1 Introduction
The analysis of volatility using high-frequency financial data is a vibrant research area in financial econometrics and statistics. In practice, understanding the dynamics of asset return volatility is essential for applications, such as hedging, option pricing, risk management, and portfolio optimization. The widespread availability of high-frequency financial data has led to the development of several effective non-parametric methods for estimating integrated volatility. Examples of these methods include two-time scale realized volatility (TSRV) (Zhang et al., 2005), multi-scale realized volatility (MSRV) (Zhang, 2006, 2011), pre-averaging realized volatility (PRV) (Christensen et al., 2010; Jacod et al., 2009), wavelet realized volatility (WRV) (Fan and Wang, 2007), kernel realized volatility (KRV) (Barndorff-Nielsen et al., 2008, 2011a), quasi-maximum likelihood estimator (QMLE) (Aït-Sahalia et al., 2010; Xiu, 2010), local method of moments (Bibinger et al., 2014), and robust pre-averaging realized volatility (Fan and Kim, 2018; Shin et al., 2023).
@@
The incorporation of high-frequency data has significantly deepened our understanding of low-frequency (e.g., daily) market dynamics. Several conditional volatility models have been proposed based on realized volatility to capture these dynamics. Examples include realized volatility-based modeling approaches (Andersen et al., 2003), heterogeneous autoregressive (HAR) models (Corsi, 2009), high-frequency-based volatility (HEAVY) models (Shephard and Sheppard, 2010), realized GARCH models (Hansen et al., 2012), and unified GARCH-Itô models (Kim and Wang, 2016; Song et al., 2021). Their empirical studies typically focus on volatility dynamics given the high-frequency information for a finite number of assets. However, in practice, we often need to manage large portfolios, which causes the overparameterization issue due to the excessive number of parameters compared to the sample size. To address this issue, approximate factor model structures are commonly imposed on large volatility matrices (Fan et al., 2013). In particular, high-dimensional factor-based Itô processes are frequently used under sparsity assumptions on the idiosyncratic volatility (Ait-Sahalia and Xiu, 2017; Fan et al., 2016a; Fan and Kim, 2018; Kim et al., 2018). Recently, Kim and Fan (2019) proposed the factor GARCH-Itô model, and Shin et al. (2021) developed the factor and idiosyncratic VAR-Itô model, both based on high-dimensional factor-based Itô processes. These models assume that the eigenvalue sequence of latent factor volatility matrices follows either a unified GARCH-Itô structure (Kim and Wang, 2016) or a VAR structure, which allows the dynamics of volatility to be explained by factors. They restrict that the eigenvectors of the latent factor volatility matrices are time-invariant. However, several empirical studies have shown that eigenvectors are time-varying (Kong and Liu, 2018; Kong et al., 2023; Su and Wang, 2017). Thus, accommodating the eigenvector dynamics is important to better account for the dynamics of large volatility matrices.
@@
This paper proposes a novel prediction approach for a future large volatility matrix. Specifically, we represent a large volatility matrix process in a cubic (order-3 tensor) form by stacking large integrated volatility matrices over time to capture interday time series dynamics. To address the high-dimensionality problem, we impose a low-rank factor structure (Chen and Fan, 2023; De Lathauwer et al., 2000; Kolda and Bader, 2009) and sparse idiosyncratic structure on the tensor. The low-rank tensor component represents a conditional expected factor volatility tensor, which follows a semiparametric factor structure (Chen et al., 2024). To implement the proposed integrated volatility matrix model, we use the Projected-PCA method suggested by Fan et al. (2016b) to estimate the time series loading matrix, which reflects the daily volatility dynamics. For example, we project the loading matrix onto a linear space spanned by past realized volatility estimators, which allows us to study the daily volatility dynamics and to predict the one-day-ahead large volatility matrix using current observed volatility information. To account for the sparse idiosyncratic volatility structure, after removing the projected factor volatility component, we adopt the principal orthogonal component thresholding (POET) (Fan et al., 2013) procedure. We refer to this method as the Projected Tensor POET (PT-POET) procedure. We then derive convergence rates for the projected integrated volatility matrix estimator and the predicted large volatility matrix using the PT-POET approach. In an empirical study, we demonstrate that the proposed PT-POET estimator performs well in out-of-sample predictions for the one-day-ahead large volatility matrix and in a minimum variance portfolio allocation analysis using high-frequency trading data.
@@
The rest of the paper is organized as follows. Section 2 establishes the model and introduces the PT-POET method to predict the conditional expected large volatility matrix. Section 3 develops an asymptotic analysis of the PT-POET estimator. The merits of the proposed method are demonstrated by a simulation study in Section 4 and by real data application on predicting the one-day-ahead volatility matrix and portfolio allocation in Section 5. Section 6 concludes the study. All proofs are presented in Appendix A.
@@ 2.1 A Model Setup
where $\mathbf{x}_{l}=(x_{i1},\dots,x_{ld})$ is observable covariates that explain the time series loading vectors, $\phi(\mathbf{x}_{l})$ is a $(Jd)\times 1$ vector of basis functions, $\mathbf{a}_{k}$ is a $(Jd)\times 1$ vector of sieve coefficients, and $R_{k}(\mathbf{x}_{l})$ is the approximation error term. In this context, for example, $\mathbf{x}_{l}$ can be the past eigenvalues in the VAR model (Shin et al., 2021) or realized largest eigenvalues of the yesterday, last week, and last month in the HAR model (Corsi, 2009). We note that (2.8) can be written as $g_{k}(\mathbf{x}_{l})=\sum_{m=1}^{d}g_{km}(x_{lm})$, where $g_{km}(x_{lm})=\sum_{j=1}^{J}b_{j,km}\phi_{j}(x_{jm})+R_{km}(x_{jm})$. Hence, the additive component of $g_{k}$ can be estimated by the sieve method. We assume that $d$ is fixed, and the number of sieve terms $J$ grows very slowly as $D\rightarrow\infty$. In a matrix form, we can write
@@
where $l-1=t_{l,0}<\cdots<t_{l,m}=l$, the microstructural noises are random variables with a mean of zero.
@@ 3 Asymptotic Properties
Assumption 3.3 is related to the accuracy of the sieve approximation and can be satisfied using a common basis such as polynomial basis or B-splines (Chen, 2007).
@@
Theorem 3.1 represents that PT-POET consistently predicts the conditional expected volatility matrix under both the max and relative Frobenius norms. As discussed in Remark 3.2, there are $m^{-1/4}$, $\sqrt{J}/P$, and $J$ terms. We note that out-of-sample predictions with any covariates $\mathbf{x}$ require additional costs such as $\sqrt{J}$ and the supremum term $\max_{j\leq J}\sup_{x}|\phi_{j}(x)|$, whereas the result in Proposition 3.1 does not require these costs since it pertains to in-sample prediction. Under the relative Frobenius norm, the additional term $s_{p}/\sqrt{p}$ arises from estimating the factor component, while the term $s_{p}\omega_{m}^{1-\eta}$ results from estimating the sparse idiosyncratic component.
@@ 4 Simulation Study
We obtained the sparse volatility component ${\cal E}=(\bm{\Sigma}_{l})_{l=1,\dots,D+1}$ as follows: let $\mathbf{d}=\mathrm{diag}(d_{1}^{2},\dots,d_{p}^{2})$, where each $\{d_{i}\}$ was generated independently from Gamma $(\alpha,\beta)$ with $\alpha=\beta=100$. We set $s=(s_{1},\dots,s_{p})^{\prime}$ to be a sparse vector, where each $s_{i}$ was drawn from $\mathcal{N}(0,1)$ with probability $\frac{0.3}{\sqrt{p}\log{p}}$, and $s_{i}=0$ otherwise. Then, we set a sparse error covariance matrix as $\bm{\Sigma}=\mathbf{d}+ss^{\prime}-\mathrm{diag}\{s_{1}^{2},\dots,s_{p}^{2}\}$, and we let $\bm{\Sigma}_{l}=\bm{\Sigma}$ for each $l$. In the simulation, we generated $\bm{\Sigma}$ until it is positive definite. We first generated high-frequency data with $m=\{250,500,2000\}$ for 200 consecutive days, and we used the subsampled log prices of the last $D$ days. We varied $D$ from 50 to 200, and the whole simulation procedure was repeated 500 times.
@@
With the aggregated volatility matrix estimates spanning $D$ days, $\widehat{{\cal Y}}=(\widehat{\bm{\Gamma}}_{l})_{l=1,\dots,D}$, we examined the out-of-sample performance of predicting the one-day-ahead aggregated volatility matrix. For comparison, the PRVM, POET, FIVAR, T-POET, and PT-POET methods were employed to predict $E(\bm{\Gamma}_{D+1}|{\cal I}_{D})$, given the past $D$ period observation. In particular, for PT-POET, we utilized the past daily, weekly, and monthly averages of the top eigenvalues for $\mathbf{X}$ with the additive polynomial basis and $J=2$. PRVM and POET represent the PRVM estimator (i.e., $\widehat{\bm{\Sigma}}_{D}$) and the POET estimator (Fan et al., 2013) based on PRVM at the $D$th day, respectively. We also employed the FIVAR method (Shin et al., 2021) to model the factor part dynamics based on the PRVM estimator. Specifically, we used the past $D$ days observations to estimate the model parameters and the previous 21 days to estimate the time-invariant eigenvectors. We do not model the idiosyncratic dynamics to compare the performance of the factor modeling. Details can be found in Shin et al. (2021). T-POET represents the $D$th day matrix estimator based on the conventional estimation procedure for the tensor observation, $\widehat{{\cal Y}}$, without using additional covariates. The integrated volatility matrix has a low-rank plus sparse structure. Hence, to estimate the idiosyncratic component of the POET, FIVAR, T-POET, and PT-POET estimators, we employed a soft thresholding scheme and used the thresholding level $\sqrt{2\log p/m^{1/2}}$ as in Kim and Fan (2019). Their idiosyncratic volatility matrix estimators are the same.
Figure 1 presents the average log Frobenius, max, Spectral, and relative Frobenius norm errors of the future volatility matrix estimators with $D=50,100,150,200$ and $m=250,500,2000$. We note that for each simulation, the target future volatility matrix is the same for each different pair of $D$ and $m$. Figure 1 shows that the PT-POET method shows the best performance. This is because PT-POET can accurately predict the future integrated volatility matrix by leveraging the HAR-based interday volatility dynamics and time-varying eigenvector in addition to eigenvalue. In addition, the matrix errors of PT-POET tend to decrease as $D$ and $m$ increase. This finding supports the theoretical results in Section 3.
@@ 5 Empirical Study
We applied the proposed PT-POET method to large volatility matrix prediction using real high-frequency trading data for 200 assets from January 2018 to December 2019 (503 trading days). We selected the top 200 large trading volume stocks among S&P 500 in the Wharton Research Data Services (WRDS) system. We used the previous tick scheme (Andersen et al., 2003; Barndorff-Nielsen et al., 2011a; Zhang, 2011) to synchronize the high-frequency data to avoid the irregular observation time error issue, and we chose 1-min log-returns.
To employ the proposed estimation procedure and other comparison methods, we need to choose the rank $r_{1}$ and $r_{2}$. We first calculated 503 daily integrated volatility matrices using the PRVM estimation method in (4.1). Then, we estimated the rank $r_{1}$ using the procedure suggested by Ait-Sahalia and Xiu (2017) as follows:
@@
To predict the conditional expected volatility matrix $E(\bm{\Gamma}_{D+1}|{\cal I}_{D})$, we employed the PT-POET, POET, FIVAR, and T-POET methods as described in Section 4. Additionally, for robustness checks, we included PT-POET2, which is based on $r_{1}=3$ and $r_{2}=2$. We also considered FIVAR-H, which incorporates a HAR structure to predict one-day-ahead leading eigenvalues instead of using a VAR structure. For PT-POETs, we utilized the ex-post daily, weakly, and monthly realized volatility based on the first eigenvalues as covariates for $\mathbf{X}$ and used the additive polynomial basis and $J=2$. We used the rolling window scheme, where the in-sample period was one of 63, 126, or 252 days. We presented the best-performing results for FIVAR, T-POET, and PT-POETs across the range of in-sample periods. Specifically, the FIVAR method, utilizing observations from the past 252 days observations to estimate the model parameters and the previous 21 days to estimate the eigenvectors, performed best. Following Shin et al. (2021), the AR lag $h=1$ was chosen based on the Bayesian information criterion (BIC). For PT-POET and T-POET, an in-sample period of 63 days yielded the best performance. We used three different out-of-sample periods: from 2019:1 to 2019:6 (period 1), from 2019:7 to 2019:12 (period 2), from 2019:1 to 2019:12 (period 3).
@@
where $T$ is the number of days in the out-of-sample period, $\widehat{\bm{\Gamma}}_{d}^{\text{POET}}$ is the POET estimator for the $d$-th day, which is a proxy of true volatility matrix, and $\widetilde{\bm{\Gamma}}_{d}$ is one of the one-day-ahead volatility matrix estimates from PRVM, POET, FIVAR, FIVAR-H, T-POET, PT-POET, and PT-POET2 for the $d$-th day of the out-of-sample period. In addition, since the true conditional expected large volatility matrix is unknown, we conducted the Diebold and Mariano (DM) test (Diebold and Mariano, 2002) using MSPE and QLIKE to assess the significance of differences in predictive performance. We compared the proposed PT-POET method with other methods. Table 1 reports the results of MSPEs and QLIKEs, and Table 2 shows the $p$-values for the DM tests. We note that the QLIKE results for the PRVM estimator are omitted, as its determinant is close to zero. The results indicate that the PT-POET estimators demonstrate the best overall performance, and PT-POET and PT-POET2 show statistically similar performances. This may be because incorporating the tensor structure and projection method, along with additional covariates such as ex-post realized volatility information, enhances prediction accuracy. We note that PT-POET does not statistically outperform FIVAR-H based on the DM test using MSPE. This outcome is due to the higher variance of prediction errors with FIVAR-H compared to FIVAR and PT-POET. That is, FIVAR-H is relatively volatile. To further evaluate the performance of the proposed method, we implemented the minimum variance portfolio allocation as described below.
@@
where $\mathbf{1}=(1,\dots,1)^{\top}\in\mathbb{R}^{p}$, the gross exposure constraint $c$ varies from 1 to 3, and $\widetilde{\bm{\Gamma}}_{d}$ is one of the one-day-ahead volatility matrix estimators obtained from PRVM, POET, FIVAR, FIVAR-H, T-POET, and PT-POETs. At the beginning of each trading day, we obtained optimal portfolios based on each estimator and held these portfolios for one day. To avoid the microstructural noise effect, we calculated the realized volatility using the 10-min portfolio log-returns. We then measured the out-of-sample risk by averaging the square root of the realized volatility for each out-of-sample period. Figure 3 illustrates the out-of-sample risks of the portfolios constructed by the PRVM, POET, FIVAR, FIVAR-H, T-POET, and PT-POET estimators. As shown in Figure 3, the PT-POET and PT-POET2 estimators demonstrate stable performance and consistently outperform the other estimators. Interestingly, while T-POET performs well in terms of global minimum risk, it becomes unstable as the gross exposure constraint increases. Additionally, PRVM and POET do not perform well. This may be because they cannot capture the dynamics of the volatility process. By considering the vector auto-regressive structure on eigenvalues of the volatility matrix, FIVAR shows improved performance compared to POET. However, FIVAR underperforms PT-POET because it does not consider the time-varying eigenvector. When comparing PT-POET and PT-POET2, PT-POET2 shows a more stable performance. This may be because adding an additional dimension of time series helps account for the eigenvector dynamics. Overall, these results indicate that the PT-POET method can efficiently predict the large integrated volatility matrix by incorporating interday time series dynamics based on both time-varying eigenvector and eigenvalue structures.

### v2

# Large Volatility Matrix Prediction using Tensor Factor Structure

@@ Abstract
Several approaches for predicting large volatility matrices have been developed based on high-dimensional factor-based Itô processes. These methods often impose restrictions to reduce the model complexity, such as constant eigenvectors or factor loadings over time. However, several studies indicate that eigenvector processes are also time-varying. To address this feature, this paper generalizes the factor structure by representing the integrated volatility matrix process as a cubic (order-3 tensor) form, which is decomposed into low-rank tensor and idiosyncratic tensor components. To predict conditional expected large volatility matrices, we propose the Projected Tensor Principal Orthogonal componEnt Thresholding (PT-POET) procedure and establish its asymptotic properties. The advantages of PT-POET are validated through a simulation study and demonstrated in an application to minimum variance portfolio allocation using high-frequency trading data.
@@ 1 Introduction
The study of volatility using high-frequency financial data is a pivotal area of research in financial econometrics and statistics. Understanding the dynamics of asset return volatility is critical for practical applications, including hedging, option pricing, risk management, and portfolio optimization. The growing availability of high-frequency financial data has spurred the development of numerous effective non-parametric methods for estimating integrated volatility. Notable examples include two-time scale realized volatility (TSRV) (Zhang et al., 2005), multi-scale realized volatility (MSRV) (Zhang, 2006, 2011), pre-averaging realized volatility (PRV) (Christensen et al., 2010; Jacod et al., 2009), wavelet realized volatility (WRV) (Fan and Wang, 2007), kernel realized volatility (KRV) (Barndorff-Nielsen et al., 2008, 2011a), quasi-maximum likelihood estimator (QMLE) (Aït-Sahalia et al., 2010; Xiu, 2010), local method of moments (Bibinger et al., 2014), and robust pre-averaging realized volatility (Fan and Kim, 2018; Shin et al., 2023).
@@
The use of high-frequency data has greatly enhanced our understanding of market dynamics at lower (e.g., daily) frequencies. To capture these dynamics, various conditional volatility models based on realized volatility have been developed. Examples include realized volatility-based modeling approaches (Andersen et al., 2003), heterogeneous autoregressive (HAR) models (Corsi, 2009), high-frequency-based volatility (HEAVY) models (Shephard and Sheppard, 2010), realized GARCH models (Hansen et al., 2012), and unified GARCH-Itô models (Kim and Wang, 2016; Song et al., 2021). Their empirical studies typically focus on volatility dynamics, given the high-frequency information for a finite number of assets. However, in practice, we often need to manage large portfolios, which causes the overparameterization issue due to the excessive number of parameters compared to the sample size. To address this issue, approximate factor model structures are commonly imposed on large volatility matrices (Fan et al., 2013). In particular, high-dimensional factor-based Itô processes are frequently used under sparsity assumptions on the idiosyncratic volatility (Ait-Sahalia and Xiu, 2017; Fan et al., 2016a; Fan and Kim, 2018; Kim et al., 2018). Recently, Kim and Fan (2019) proposed the factor GARCH-Itô model, and Shin et al. (2021) developed the factor and idiosyncratic VAR-Itô model, both based on high-dimensional factor-based Itô processes. These models assume that the eigenvalue sequence of latent factor volatility matrices follows either a unified GARCH-Itô structure (Kim and Wang, 2016) or a VAR structure, which allows the dynamics of volatility to be explained by factors. They restrict that the eigenvectors of the latent factor volatility matrices are time-invariant. However, several empirical studies have shown that eigenvectors are time-varying (Kong and Liu, 2018; Kong et al., 2023; Su and Wang, 2017). Thus, accommodating the eigenvector dynamics is essential to better account for the dynamics of large volatility matrices.
@@
This paper proposes a novel prediction approach for a future large volatility matrix. Specifically, we represent a large volatility matrix process in a cubic (order-3 tensor) form by stacking large integrated volatility matrices over time to capture interday time series dynamics. To address the high-dimensionality problem, we impose a low-rank factor structure (Chen and Fan, 2023; De Lathauwer et al., 2000; Kolda and Bader, 2009) and sparse idiosyncratic structure on the tensor. The low-rank tensor component represents a conditional expected factor volatility tensor, which follows a semiparametric factor structure (Chen et al., 2024), and we apply the Projected-PCA (Fan et al., 2016b) procedure to estimate the time series loading matrix. To account for the sparse idiosyncratic volatility structure, after removing the projected factor volatility component, we adopt the principal orthogonal component thresholding (POET) (Fan et al., 2013) procedure. This method is called the Projected Tensor POET (PT-POET) procedure. We then derive convergence rates for the projected integrated volatility matrix estimator and the predicted large volatility matrix using the PT-POET approach. In an empirical study, we demonstrate that the proposed PT-POET estimator performs well in out-of-sample predictions for the one-day-ahead large volatility matrix and in a minimum variance portfolio allocation analysis using high-frequency trading data.
@@
The rest of the paper is organized as follows. Section 2 establishes the model and introduces the PT-POET method to predict the conditional expected large volatility matrix. Section 3 develops an asymptotic analysis of the PT-POET estimator. The merits of the proposed method are demonstrated by a simulation study in Section 4 and by real data application on predicting the one-day-ahead volatility matrix and portfolio allocation in Section 5. Section 6 concludes the study. All proofs are presented in the online supplement file.
@@ 2.1 A Model Setup
where $\mathbf{x}_{l}=(x_{i1},\dots,x_{ld})$ is observable covariates that explain the time series loading vectors, $\phi(\mathbf{x}_{l})$ is a $(Jd)\times 1$ vector of basis functions, $\mathbf{a}_{k}$ is a $(Jd)\times 1$ vector of sieve coefficients, and $R_{k}(\mathbf{x}_{l})$ is the approximation error term. In this context, for example, $\mathbf{x}_{l}$ can be the past eigenvalues in the VAR model (Shin et al., 2021) or realized largest eigenvalues of yesterday, last week, and last month in the HAR model (Corsi, 2009). We note that (2.8) can be written as $g_{k}(\mathbf{x}_{l})=\sum_{m=1}^{d}g_{km}(x_{lm})$, where $g_{km}(x_{lm})=\sum_{j=1}^{J}b_{j,km}\phi_{j}(x_{jm})+R_{km}(x_{jm})$. Hence, the additive component of $g_{k}$ can be estimated by the sieve method. We assume that $d$ is fixed, and the number of sieve terms $J$ grows very slowly as $D\rightarrow\infty$. In a matrix form, we can write
@@
where $l-1=t_{l,0}<\cdots<t_{l,m}=l$, and the microstructural noises are random variables with a mean of zero.
@@ 3 Asymptotic Properties
Assumption 3.3 is related to the accuracy of the sieve approximation and can be satisfied using a common basis such as a polynomial basis or B-splines (Chen, 2007).
@@
Theorem 3.1 shows that PT-POET consistently predicts the conditional expected volatility matrix under both the max and relative Frobenius norms. As discussed in Remark 3.2, there are $m^{-1/4}$, $\sqrt{J}/P$, and $J$ terms. We note that out-of-sample predictions with any covariates $\mathbf{x}$ require additional costs such as $\sqrt{J}$ and the supremum term $\max_{j\leq J}\sup_{x}|\phi_{j}(x)|$, whereas the result in Proposition 3.1 does not require these costs since it pertains to in-sample prediction. Under the relative Frobenius norm, the additional term $s_{p}/\sqrt{p}$ arises from estimating the factor component, while the term $s_{p}\omega_{m}^{1-\eta}$ results from estimating the sparse idiosyncratic component.
@@ 4 Simulation Study
We obtained the sparse volatility component ${\cal E}=(\bm{\Sigma}_{l})_{l=1,\dots,D+1}$ as follows: let $\mathbf{d}=\mathrm{diag}(d_{1}^{2},\dots,d_{p}^{2})$, where each $\{d_{i}\}$ was generated independently from Gamma $(\alpha,\beta)$ with $\alpha=\beta=100$. We set $s=(s_{1},\dots,s_{p})^{\prime}$ to be a sparse vector, where each $s_{i}$ was drawn from $\mathcal{N}(0,1)$ with probability $\frac{0.3}{\sqrt{p}\log{p}}$, and $s_{i}=0$ otherwise. Then, we set a sparse error covariance matrix as $\bm{\Sigma}=\mathbf{d}+ss^{\prime}-\mathrm{diag}\{s_{1}^{2},\dots,s_{p}^{2}\}$, and we let $\bm{\Sigma}_{l}=\bm{\Sigma}$ for each $l$. In the simulation, we generated $\bm{\Sigma}$ until it is positive definite. We first generated high-frequency data with $m=\{250,500,2000\}$ for 200 consecutive days and used the subsampled log prices of the last $D$ days. We varied $D$ from 50 to 200, and the whole simulation procedure was repeated 500 times.
@@
With the aggregated volatility matrix estimates spanning $D$ days, $\widehat{{\cal Y}}=(\widehat{\bm{\Gamma}}_{l})_{l=1,\dots,D}$, we examined the out-of-sample performance of predicting the one-day-ahead aggregated volatility matrix. For comparison, the PRVM, POET, FIVAR, T-POET, and PT-POET methods were employed to predict $E(\bm{\Gamma}_{D+1}|{\cal I}_{D})$, given the past $D$ period observation. In particular, for PT-POET, we utilized the past daily, weekly, and monthly averages of the top eigenvalues for $\mathbf{X}$ with the additive polynomial basis and $J=2$. PRVM and POET represent the PRVM estimator (i.e., $\widehat{\bm{\Sigma}}_{D}$) and the POET estimator (Fan et al., 2013) based on PRVM at the $D$th day, respectively. We also employed the FIVAR method (Shin et al., 2021) to model the factor part dynamics based on the PRVM estimator. Specifically, we used the past $D$ days’ observations to estimate the model parameters and the previous 21 days to estimate the time-invariant eigenvectors. We do not model the idiosyncratic dynamics to compare the performance of the factor modeling. Details can be found in Shin et al. (2021). T-POET represents the $D$th day matrix estimator based on the conventional estimation procedure for the tensor observation, $\widehat{{\cal Y}}$, without using additional covariates. The integrated volatility matrix has a low-rank plus sparse structure. Hence, to estimate the idiosyncratic component of the POET, FIVAR, T-POET, and PT-POET estimators, we employed a soft thresholding scheme and used the thresholding level $\sqrt{2\log p/m^{1/2}}$ as in Kim and Fan (2019). Their idiosyncratic volatility matrix estimators are the same.
Figure 1 presents the average log Frobenius, max, Spectral, and relative Frobenius norm errors of the future volatility matrix estimators with $D=50,100,150,200$ and $m=250,500,2000$. We note that for each simulation, the target future volatility matrix is the same for each different pair of $D$ and $m$. Figure 1 shows that the PT-POET method performs best. This is because PT-POET can accurately predict the future integrated volatility matrix by leveraging the HAR-based interday volatility dynamics and time-varying eigenvector in addition to eigenvalue. In addition, the matrix errors of PT-POET tend to decrease as $D$ and $m$ increase. This finding supports the theoretical results in Section 3.
@@ 5 Empirical Study
We applied the proposed PT-POET method to large volatility matrix prediction using real high-frequency trading data for 200 assets from January 2018 to December 2019 (503 trading days). We selected the top 200 large trading volume stocks in the S&P 500 in the Wharton Research Data Services (WRDS) system. We used the previous tick scheme (Andersen et al., 2003; Barndorff-Nielsen et al., 2011a; Zhang, 2011) to synchronize the high-frequency data to avoid the irregular observation time error issue, and we chose 1-min log-returns.
We need to choose the ranks $r_{1}$ and $r_{2}$ to employ the proposed estimation procedure and other comparison methods. We first calculated 503 daily integrated volatility matrices using the PRVM estimation method in (4.1). Then, we estimated the rank $r_{1}$ using the procedure suggested by Ait-Sahalia and Xiu (2017) as follows:
@@
To predict the conditional expected volatility matrix $E(\bm{\Gamma}_{D+1}|{\cal I}_{D})$, we employed the PT-POET, POET, FIVAR, and T-POET methods as described in Section 4. Additionally, for robustness checks, we included PT-POET2, which is based on $r_{1}=3$ and $r_{2}=2$. We also considered FIVAR-H, which incorporates a HAR structure instead of a VAR structure to predict one-day-ahead leading eigenvalues. For PT-POETs, we utilized the ex-post daily, weakly, and monthly realized volatility based on the first eigenvalues as covariates for $\mathbf{X}$ and used the additive polynomial basis and $J=2$. We used the rolling window scheme, where the in-sample period was one of 63, 126, or 252 days. We presented the best-performing results for FIVAR, T-POET, and PT-POETs across the range of in-sample periods. Specifically, the FIVAR method, utilizing observations from the past 252 days observations to estimate the model parameters and the previous 21 days to estimate the eigenvectors, performed best. Following Shin et al. (2021), the AR lag $h=1$ was chosen based on the Bayesian information criterion (BIC). For PT-POET and T-POET, an in-sample period of 63 days yielded the best performance. We used three different out-of-sample periods: from 2019:1 to 2019:6 (period 1), from 2019:7 to 2019:12 (period 2), and from 2019:1 to 2019:12 (period 3).
@@
where $T$ is the number of days in the out-of-sample period, $\widehat{\bm{\Gamma}}_{d}^{\text{POET}}$ is the POET estimator for the $d$-th day, which is a proxy of true volatility matrix, and $\widetilde{\bm{\Gamma}}_{d}$ is one of the one-day-ahead volatility matrix estimates from PRVM, POET, FIVAR, FIVAR-H, T-POET, PT-POET, and PT-POET2 for the $d$-th day of the out-of-sample period. In addition, since the true conditional expected large volatility matrix is unknown, we conducted the Diebold and Mariano (DM) test (Diebold and Mariano, 1995) using MSPE and QLIKE to assess the significance of differences in predictive performance. We compared the proposed PT-POET method with other methods. Table 1 reports the results of MSPEs and QLIKEs, and Table 2 shows the $p$-values for the DM tests. We note that the QLIKE results for the PRVM estimator are omitted, as its determinant is close to zero. The results indicate that the PT-POET estimators demonstrate the best overall performance, and PT-POET and PT-POET2 show statistically similar performances. This may be because incorporating the tensor structure and projection method, along with additional covariates such as ex-post realized volatility information, enhances prediction accuracy. We note that PT-POET does not statistically outperform FIVAR-H based on the DM test using MSPE. This outcome is due to the higher variance of prediction errors with FIVAR-H compared to FIVAR and PT-POET. That is, FIVAR-H is relatively volatile. To further evaluate the proposed method’s performance, we implemented the minimum variance portfolio allocation, as described below.
@@
where $\mathbf{1}=(1,\dots,1)^{\top}\in\mathbb{R}^{p}$, the gross exposure constraint $c$ varies from 1 to 3, and $\widetilde{\bm{\Gamma}}_{d}$ is one of the one-day-ahead volatility matrix estimators obtained from PRVM, POET, FIVAR, FIVAR-H, T-POET, and PT-POETs. At the beginning of each trading day, we obtained optimal portfolios based on each estimator and held these portfolios for one day. We calculated the realized volatility using the 10-min portfolio log-returns to avoid the microstructural noise effect. We then measured the out-of-sample risk by averaging the square root of the realized volatility for each out-of-sample period. Figure 3 illustrates the out-of-sample risks of the portfolios constructed by the PRVM, POET, FIVAR, FIVAR-H, T-POET, and PT-POET estimators. As shown in Figure 3, the PT-POET and PT-POET2 estimators demonstrate stable performance and consistently outperform the other estimators. Interestingly, while T-POET performs well in terms of global minimum risk, it becomes unstable as the gross exposure constraint increases. Additionally, PRVM and POET do not perform well. This may be because they cannot capture the dynamics of the volatility process. By considering the vector auto-regressive structure on eigenvalues of the volatility matrix, FIVAR shows improved performance compared to POET. However, FIVAR underperforms PT-POET because it does not consider the time-varying eigenvector. When comparing PT-POET and PT-POET2, PT-POET2 shows a more stable performance. This may be because adding an additional dimension of time series helps account for the eigenvector dynamics. Overall, these results indicate that the PT-POET method can efficiently predict the large integrated volatility matrix by incorporating interday time series dynamics based on both time-varying eigenvector and eigenvalue structures.